OpenStudy (anonymous):

integrate sqrt(1+e^(2x))..help please

hartnn (hartnn):

did you try the u sub \(\Large u= e^x\) ?

hartnn (hartnn):

a trig sub will be better \(\Large e^x = \tan u\)

OpenStudy (anonymous):

i will try it

hartnn (hartnn):

what will be dx = ....? in terms of 'u' only?

ganeshie8 (ganeshie8):

Alternatively if you like partial fractions, try \(\sqrt{1+e^{2x}} = u \) it gives a nice fraction to integrate :)

OpenStudy (michele_laino):

please try this substitution: \[\sqrt{1+e ^{2x}}=e ^{x}+t\] where t is your new variable of integration. Please square both sides and try to solve for e^x you should get this: \[e ^{x}=\frac{ 1-t ^{2} }{ 2t }\]

hartnn (hartnn):

Awesome! so many different ways to integrate the same function :D